OpenStudy (anonymous):

How do I solve this? cos2X=-cosX on [0, 2pi)

OpenStudy (fifciol):

use this : \[\cos2x=\cos^2x-\sin^2x\]

OpenStudy (fifciol):

and this: \[\sin^2x + \cos^2x=1\]

OpenStudy (fifciol):

so you have then:\[\cos2x=\cos^2x-(1-\cos^x)=2\cos^2x-1\]

OpenStudy (fifciol):

substitute that in initial equation, sypstitute t= cosx and you will have quadratic equation. Solve for t and then for x

OpenStudy (fifciol):

choose only the x's from interval [0, 2pi)

OpenStudy (fifciol):

Do you know how to do that?

OpenStudy (anonymous):

Yes. I understand how you came to 2cosx^2 -1 but I do not understand after that.

OpenStudy (fifciol):

If you substitute that in your initial equation you have: \[2\cos^2x+cosx-1=0\]

OpenStudy (fifciol):

t=cosx and \[t \in<-1;1>\] so \[2t^2 + t -1=0\]

OpenStudy (fifciol):

Do you understand why t must be between -1 and 1?

OpenStudy (fifciol):